# Deepfake Generation and Detection: A Benchmark and Survey

### v1

@@ Abstract
Deepfake is a technology dedicated to creating highly realistic facial images and videos under specific conditions, such as identity-aware face swapping and audio-guided talking face generation. This technology has significant application potential in fields such as entertainment, movie production, digital human creation, to name a few. With the advancements in deep learning, techniques primarily represented by Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) have achieved impressive generation results. More recently, the emergence of diffusion models with powerful image and generation capabilities has sparked a renewed wave of research and industrial interest in this technology. In addition to the advancements in deepfake generation, corresponding detection technologies need to continuously evolve to regulate the potential misuse of deepfakes, such as for privacy invasion and phishing attacks. This survey comprehensively reviews the latest developments in deepfake generation and detection, summarizing and analyzing the current state of the art in this rapidly evolving field. We first unify task definitions, comprehensively introduce datasets and metrics, and discuss the development of generation and detection technology frameworks. Then, we discuss the development of several related sub-fields and focus on researching four mainstream deepfake fields: popular face swap, face reenactment, talking face generation, and facial attribute editing, as well as foreign detection. Subsequently, we comprehensively benchmark representative methods on popular datasets for each field, fully evaluating the latest and influential works published in top conferences/journals. Finally, we analyze the challenges and future research directions of the discussed fields. We closely follow the latest developments in this project.
@@
Deepfake Generation, Face Swapping, Face Reenactment, Talking Face Generation, Facial Attribute Editing, Foreign Detection, Survey
@@ 1 Introduction
Figure 1: A time diagram that reflects the survey pipeline. Zoom in for a better holistic perception of this work.
While enjoying the novelty and convenience of this technology, the unethical use of it raises concerns over the spread of privacy invasion, dissemination of fake news, and phishing attacks, necessitating the development of effective foreign detection methods [30] in opposition, a.k.a., deepfake detection. From the earliest handcrafted feature-based methods [31, 32] to deep learning-based methods [33, 34], and the recent hybrid detection techniques [35], forgery detection has undergone substantial technological advancements along with the development of generative technologies. The data modality has also transitioned from the spatial and frequent domains [36, 37] to the more challenging temporal domain [38, 39]. Considering that current generative technologies have a significantly higher level of interest, develop faster, and can generate indistinguishable content from reality [40], corresponding detection technologies need continuous evolving.
@@
$\bullet$ Contribution. In this survey, we comprehensively explore the key technologies and latest advancements in Deepfakes generation and foreign detection. We first unify the task definitions (Sec. 2.1), provide a comprehensive comparison of datasets and metrics (Sec. 2.3), and discuss the development of related technologies. Specifically, we investigate four mainstream deepfake fields: face swapping (Sec. 3.1.1), face reenactment (Sec. 3.1.2), talking face generation (Sec. 3.1.3), and facial attribute editing (mainly on multiple editing) (Sec. 3.1.4), as well as foreign detection (Sec. 3.2). We also analyze the benchmarks and settings for each domain, thoroughly evaluating the latest and influential works published in top-tier conferences/journals (Sec. 4), especially recent diffusion-based approaches. Additionally, we discuss closely related fields, including head swap, face super-resolution, face reconstruction, face inpainting, body animation, portrait style transfer, makeup transfer, and adversarial sample detection. Benefiting from the current popularity of AIGC, the research iteration cycle in the deepfake field has been significantly reduced, and we keep updating and discussing the in-submission works in the revised version.
$\bullet$ Scope. This survey primarily focuses on mainstream face-related tasks, including face swapping, face reenactment, talking face generation, facial (multiple) attribute editing, and foreign detection. We also cover some related domain tasks in Sec. 2.4 and detail specific popular sub-tasks in Sec. 3.3. Considering the large number of articles (including published and preprints), we mainly include representative and attention-grabbing works. In addition, we compare this investigation with recent surveys. Sha et al[1] only discuss character generation while we cover a more comprehensive range of tasks. Compared to works [40, 42, 43], our study encompasses a broader range of technical models, particularly the more powerful diffusion-based methods. Additionally, we thoroughly discuss the related sub-fields of deepfake generation and detection.
$\bullet$ Survey Pipeline. Fig. 1 shows the pipeline of this survey. Sec. 2 provides the essential background knowledge that encompasses task-specific definitions, datasets, evaluation metrics, and exploration of related research areas. Sec. 3 presents technical discussions, examining four most popular deepfake tasks from the perspective of technological categorization and evolution. In addition, we meticulously classify and discuss foreign detection techniques, emphasizing the technological routes involved. Subsequently, Sec. 4 organizes and evaluates performance from various approaches for a thorough and fair performance comparison. Sec. 5 focuses on a critical review of the challenges that persist in existing techniques, outlining feasible future development directions. Finally, we encapsulate the entirety of the paper through a comprehensive summary in Sec. 6.
@@ 2.1 Problem Definition
$\bullet$ Foreign Detection aims to detect and identify anomalies, tampering, or forgery areas in images or videos by the anomaly score $S_{o}$, and it has great research and application value in information security and multimedia forensics.
@@ 2.2 Technological History and Roadmap
$\bullet$ Discriminative Neural Network. Convolutional Neural Networks (CNNs) [51, 52, 53, 54, 55, 56] have played a pivotal role in the history of deep learning, undergoing development over nearly 40 years. LeNet [51], as the pioneer of CNNs, showcased the charm of machine learning. AlexNet [52] and ResNet [53] made deep learning feasible. Recently, ConvNeXt [54] has achieved excellent results surpassing those of Swin-Transformer [57]. The Transformer architecture initially proposed [58] in 2017. The core idea involves using self-attention mechanisms to capture dependencies between different positions in the input sequence, enabling global modeling of sequences. ViT [59] demonstrates that using Transformer in the field of computer vision can still achieve excellent performance and PVT [60] overcomes the challenges of adapting Transformer to various dense prediction tasks. In addition, Swin-Transformer [57] addresses the limitations of Transformer in handling high-resolution image processing tasks. Recently, Swin-Transformer V2 [61] further improves the model’s efficiency and the resolution of manageable inputs.
@@
$\bullet$ Work Summary. The evolution of mainstream generative models is depicted chronologically in Fig. 3. This survey delves into four categories of generation tasks along with the foreign detection task, and the publication years distribution of the surveyed articles is shown in Fig. 4.
@@
Foreign Detection DeepfakeTIMIT [77] Audio-Video 640 videos The dataset is evenly divided into two versions: LQ (64$\times$64) and HQ (128$\times$128), with all videos using face swapping forgery.
@@
2.3 Datasets, Metrics, and Loss Function
$\bullet$ Dataset. Given the various datasets in surveyed fields, we use numerical labels to save post-textual space. 1) Commonly used deepfake generation datasets include LFW [68], CelebA [69], CelebA-HQ [72], VGGFace [70], VGGFace2 [73], FFHQ [5], Multi-PIE [85], VoxCeleb1 [71], VoxCeleb2 [74], MEAD [75], MM CelebA-HQ [86], CelebAText-HQ [87], CelebV-HQ [76], TalkingHead-1KH [88], LRS2 [89], LRS3 [90], etc. 2) Commonly used foreign detection datasets include UADFV [91], DeepfakeTIMIT [77], FF++ [78], Deeperforensics-1.0 [81], DFDCP [79], DFDC [82], Celeb-DF [83], Celeb-DFv2 [83], FakeAVCeleb [84], DFD [80], WildDeepfake [92], KoDF [93], UADFV [94], etc. We summarize popular datasets in Tab. I.
Figure 4: Works summarization on different directions per year. Data is obtained on 2024/03/26.
$\bullet$ Metric. 1) For deepfake generation tasks, commonly used metrics include: Peak Signal-to-Noise Ratio (PSNR) [95], Structured Similarity (SSIM) [96], Learned Perceptual Image Patch Similarity (LPIPS) [97], Fréchet Inception Distance (FID) [98], Kernel Inception Distance (KID) [99], Cosine Similarity (CSIM) [88], Identity Retrieval Rate (ID Ret) [100], Expression Error [101], Pose Error [102], Landmark Distance (LMD) around the mouths [103], Lip-sync Confidence (LSE-C) [104], Lip-sync Distance (LSE-D) [104], etc. 2) For foreign generation commonly uses: Area Under the ROC Curve (AUC) [105], Accuracy (ACC) [106], Equal Error Rate (EER) [107], Average Precision (AP) [30], F1-Score [108], etc. Detailed definitions will be thoroughly explained in Appendix A.
@@ 2.4 Related Research Domains
3 Methods: A Survey
This section systematically examines four generation tasks:Face Swapping (Sec. 3.1.1), Talking Face Generation (Sec. 3.1.3), Face Reenactment (Sec. 3.1.2), Facial Attribute Editing (Sec. 3.1.4). Additionally, we review techniques for foreign detection (commonly known as deepfake detection) in Sec. 3.2. We document and compare the most representative works in the main text. Furthermore, we discuss several related domains that garnered significant attention in Sec. 3.3.
@@ 3.1 Deepfake Generation
S2Swap [189] MM’23 ➊➋➌➑ G.+3D Poor ability to preserve face feature attributes. Achieving high-fidelity face swapping through semantic disentanglement and structural enhancement.
@@
Other Cui et al[193] CVPR’23 ➊➋ Other Altered target image lighting details. Introducing a multiscale transformer network focusing on high-quality semantically aware correspond- ences between source and target faces.
@@ 3.1.1 Face Swapping
4) Geometric Detail Preservation. To capture and reproduce more facial geometric details, some methods [204, 187, 188, 205] introduce 3DMM into GANs, enabling the incorporation of 3D priors. HifiFace [204] introduces a novel 3D shape-aware identity extractor, replacing traditional face recognition networks to generate identity vectors that include precise shape information. FlowFace [188] introduces a two-stage framework based on semantic guidance to achieve shape-aware face swapping. FlowFace++ [205] improves upon FlowFace by utilizing a pre-trained Mask Autoencoder (MAE) to convert face images into a fine-grained representation space shared between the target and source faces. It further enhances feature fusion for both source and target by introducing a cross-attention fusion module. However, most of the aforementioned methods often struggle to effectively handle occlusion issues.
@@
$\bullet$ Other Techniques. There are also some methods that stand independently from the above classifications that are discussed here collectively. Fast Face-swap [211] views the identity swap task as a style transfer task, achieving its goals based on VGG-Net. However, this method has poor generalization. Some methods [193, 194] apply the Transformer architecture to face swapping tasks. Leveraging a facial encoder based on the Swin Transformer [57], TransFS [194] obtains rich facial features, enabling facial swapping in high-resolution images. ReliableSwap [195] enhances the model’s identity preservation capabilities by constructing a reliable supervisor called the "cyclic triplet." However, it has limitations in preserving attribute information.
@@ 3.1.2 Face Reenactment
$\bullet$ Face Feature Decoupling. The latent feature decoupling and driving methods [16, 228, 227, 232, 233] aims to disentangle facial features in the latent space of the driving video, replacing or mapping the corresponding latent information to achieve high-fidelity facial reproduction under specific conditions. HyperReenact [16] uses attribute decoupling, employing a hyper-network to refine source identity features and modify facial poses. StyleMask [228] separates facial pose and expression from the identity information of the source image by learning masks and blending corresponding channels in the pre-trained style space S of StyleGAN2. HiDe-NeRF [227] employs a deformable neural radiance field to represent a 3D scene, with a lightweight deformation module explicitly decoupling facial pose and expression attributes.
$\bullet$ Self-supervised Method. Self-supervised learning employs supervisory signals inferred from the intrinsic structure of the data, reducing the reliance on external data labels [229, 234, 230, 235]. Oorloff et al[230] employs self-supervised methods to train an encoder, disentangling identity and facial attribute information of portrait images within the pre-defined latent space itself of a pre-trained StyleGAN2. Zhang et al[235] utilizes 3DMM to provide geometric guidance, employs pre-computed optical flow to guide the estimation of motion fields, and relies on pre-computed occlusion maps to guide the perception and repair of occluded areas.
TABLE IV: Overview of representative talking face generation methods. Notations: ➊ LRW, ➋ VoxCeleb2, ➌ MEAD, ➍ Self-build, ➎ LRS2, ➏ HDTF, ➐ LRS3, ➑ CREMA-D, ➒ VoxCeleb, ➓ FFHQ.
@@
AD-NeRF [236] ICCV’21 Inadequate control of emotions and latent attributes. The NeRF based approach achieves accurate reproduction of detailed facial components and generates the upper body region.
SPACE [238] ICCV’23 ➋➌ Lack emotional and other latent attributes control Constructed a novel facial intermediate representation, achieving control overhead pose, blinking, and gaze direction.

### v3

@@ Abstract
Deepfake is a technology dedicated to creating highly realistic facial images and videos under specific conditions, which has significant application potential in fields such as entertainment, movie production, digital human creation, to name a few. With the advancements in deep learning, techniques primarily represented by Variational Autoencoders and Generative Adversarial Networks have achieved impressive generation results. More recently, the emergence of diffusion models with powerful generation capabilities has sparked a renewed wave of research. In addition to deepfake generation, corresponding detection technologies continuously evolve to regulate the potential misuse of deepfakes, such as for privacy invasion and phishing attacks. This survey comprehensively reviews the latest developments in deepfake generation and detection, summarizing and analyzing current state-of-the-arts in this rapidly evolving field. We first unify task definitions, comprehensively introduce datasets and metrics, and discuss developing technologies. Then, we discuss the development of several related sub-fields and focus on researching four representative deepfake fields: face swapping, face reenactment, talking face generation, and facial attribute editing, as well as forgery detection. Subsequently, we comprehensively benchmark representative methods on popular datasets for each field, fully evaluating the latest and influential published works. Finally, we analyze challenges and future research directions of the discussed fields.
@@
Deepfake Generation, Face Swapping, Face Reenactment, Talking Face Generation, Facial Attribute Editing, Forgery detection, Survey
@@ 1 Introduction
Figure 1: Time diagram that reflects the survey pipeline. Zoom in for a better holistic perception of this work.
While enjoying the novelty and convenience of this technology, the unethical use of it raises concerns over the spread of privacy invasion, and dissemination of fake news, necessitating the development of effective forgery detection methods [30] in opposition, a.k.a., deepfake detection. From the earliest handcrafted feature-based methods [31, 32] to deep learning-based methods [33, 34], and the recent hybrid detection techniques [35], forgery detection has undergone substantial technological advancements along with the development of generative technologies. The data modality has also transitioned from the spatial and frequent domains [36, 37] to the more challenging temporal domain [38, 39]. Considering that current generative technologies have a higher level of interest, develop faster, and can generate indistinguishable content from reality [40], corresponding detection technologies need continuous evolution.
@@
$\bullet$ Contribution. In this survey, we comprehensively explore the key technologies and latest advancements in Deepfakes generation and forgery detection. We first unify the task definitions (Sec. 2.1), provide a comprehensive comparison of datasets and metrics (Sec. 2.3), and discuss the development of related technologies. Specifically, we investigate four mainstream deepfake fields: face swapping (Sec. 3.1.1), face reenactment (Sec. 3.1.2), talking face generation (Sec. 3.1.3), and facial attribute editing (mainly on multiple editing) (Sec. 3.1.4), as well as forgery detection (Sec. 3.2). We also analyze the benchmarks and settings for each domain, thoroughly evaluating the latest and influential works published in top-tier conferences/journals (Sec. 4), especially recent diffusion-based approaches. Additionally, we discuss closely related fields, including head swapping, face super-resolution, face reconstruction, face inpainting, body animation, portrait style transfer, makeup transfer, and adversarial sample detection. Benefiting from the current popularity of AIGC, the research iteration cycle in the deepfake field has been significantly reduced, and we keep updating and discussing the in-submission works in the revised version.
$\bullet$ Scope. This survey primarily focuses on mainstream face-related tasks, including face swapping, face reenactment, talking face generation, facial (multiple) attribute editing, and forgery detection. We also cover some related domain tasks in Sec. 2.4 and detail specific popular sub-tasks in Sec. 3.3. Considering the large number of articles (including published and preprints), we mainly include representative and attention-grabbing works. In addition, we compare this investigation with recent surveys. Sha et al[1] only discuss character generation while we cover a more comprehensive range of tasks. Compared to works [40, 42, 43], our study encompasses a broader range of technical models, particularly the more powerful diffusion-based methods. Additionally, we thoroughly discuss the related sub-fields of deepfake generation and detection.
$\bullet$ Survey Pipeline. Fig. 1 shows the pipeline of this survey. Sec. 2 provides the essential background knowledge that encompasses task-specific definitions, datasets, evaluation metrics, and exploration of related research areas. Sec. 3 presents technical discussions, examining four most popular deepfake tasks from the perspective of technological categorization and evolution. In addition, we meticulously classify and discuss forgery detection techniques, emphasizing the technological routes involved. Subsequently, Sec. 4 organizes and evaluates performance from various approaches for a thorough and fair performance comparison. Sec. 5 focuses on a critical review of the challenges that persist in existing techniques, outlining feasible future development directions. Finally, we encapsulate the entirety of the paper through a comprehensive summary in Sec. 6.
@@ 2.1 Problem Definition
$\bullet$ Forgery Detection aims to detect and identify anomalies, tampering, or forgery areas in images or videos by the anomaly score $S_{o}$, and it has great research and application value in information security and multimedia forensics.
@@ 2.2 Technological History and Roadmap
$\bullet$ Discriminative Neural Network. Convolutional Neural Networks (CNNs) [51, 52, 53, 54, 55, 56] have played a pivotal role in the history of deep learning. LeNet [51], as the pioneer of CNNs, showcased the charm of machine learning. AlexNet [52] and ResNet [53] made deep learning feasible. Recently, ConvNeXt [54] has achieved excellent results surpassing those of Swin-Transformer [57]. The Transformer architecture initially proposed [58] in 2017. The core idea involves using self-attention mechanisms to capture dependencies between different positions in the input sequence, enabling global modeling of sequences. ViT [59] demonstrates that using Transformer in the field of computer vision can still achieve excellent performance and PVT [60] overcomes the challenges of adapting Transformer to various dense prediction tasks. In addition, Swin-Transformer [57] addresses the limitations of Transformer in handling high-resolution image processing tasks. Subsequently, Swin-Transformer V2 [61] further improves the model’s efficiency and the resolution of manageable inputs.
@@
$\bullet$ Work Summary. The evolution of mainstream generative models is depicted chronologically in Fig. 3. This survey delves into four categories of generation tasks along with the forgery detection task, and the publication years distribution of the surveyed articles is shown in Fig. 4.
@@
Forgery Detection DeepfakeTIMIT [77] Audio-Video 640 videos The dataset is evenly divided into two versions: LQ (64$\times$64) and HQ (128$\times$128), with all videos using face swapping forgery.
@@
2.3 Datasets, Metrics, and Losses
$\bullet$ Dataset. Given the various datasets in surveyed fields, we use numerical labels to save post-textual space. 1) Commonly used deepfake generation datasets include LFW [68], CelebA [69], CelebA-HQ [72], VGGFace [70], VGGFace2 [73], FFHQ [5], Multi-PIE [85], VoxCeleb1 [71], VoxCeleb2 [74], MEAD [75], MM CelebA-HQ [86], CelebAText-HQ [87], CelebV-HQ [76], TalkingHead-1KH [88], LRS2 [89], LRS3 [90], etc. 2) Commonly used forgery detection datasets include UADFV [91], DeepfakeTIMIT [77], FF++ [78], Deeperforensics-1.0 [81], DFDCP [79], DFDC [82], Celeb-DF [83], Celeb-DFv2 [83], FakeAVCeleb [84], DFD [80], WildDeepfake [92], KoDF [93], UADFV [94], Deephy [95], DF-Platter [96], etc. We summarize popular datasets in Tab. I.
Figure 4: Works summarization on different directions per year. Data is obtained on 2024/04/20.
$\bullet$ Metric. 1) For deepfake generation tasks, commonly used metrics include: Peak Signal-to-Noise Ratio (PSNR) [97], Structured Similarity (SSIM) [98], Learned Perceptual Image Patch Similarity (LPIPS) [99], Fréchet Inception Distance (FID) [100], Kernel Inception Distance (KID) [101], Cosine Similarity (CSIM) [88], Identity Retrieval Rate (ID Ret) [102], Expression Error [103], Pose Error [104], Landmark Distance (LMD) around the mouths [105], Lip-sync Confidence (LSE-C) [106], Lip-sync Distance (LSE-D) [106], etc. 2) For forgery generation commonly uses: Area Under the ROC Curve (AUC) [107], Accuracy (ACC) [108], Equal Error Rate (EER) [109], Average Precision (AP) [30], F1-Score [110], etc. Detailed definitions are explained in Appendix A.
@@ 2.4 Related Research Domains
3 Deepfake Inspections: A Survey
This section systematically examines four generation tasks:Face Swapping (Sec. 3.1.1), Talking Face Generation (Sec. 3.1.3), Face Reenactment (Sec. 3.1.2), Facial Attribute Editing (Sec. 3.1.4). Additionally, we review techniques for forgery detection (commonly known as deepfake detection) in Sec. 3.2. We document and compare the most representative works in the main text. Furthermore, we discuss several related domains that garnered significant attention in Sec. 3.3.
@@ 3.1 Deepfake Generation
Alternative Cui et al[193] CVPR’23 ➊➋ Other Altered target image lighting details. Introducing a multiscale transformer network focusing on high-quality semantically aware correspond- ences between source and target faces.
@@ 3.1.1 Face Swapping
4) Geometric Detail Preservation. To capture and reproduce more facial geometric details, some methods [203, 204, 189, 205] introduce 3DMM into GANs, enabling the incorporation of 3D priors. HifiFace [203] introduces a novel 3D shape-aware identity extractor, replacing traditional face recognition networks to generate identity vectors that include precise shape information. FlowFace [189] introduces a two-stage framework based on semantic guidance to achieve shape-aware face swapping. FlowFace++ [205] improves upon FlowFace by utilizing a pre-trained Mask Autoencoder to convert face images into a fine-grained representation space shared between the target and source faces. It further enhances feature fusion for both source and target by introducing a cross-attention fusion module. However, most of the aforementioned methods often struggle to effectively handle occlusion issues.
@@
$\bullet$ Alternative Techniques. There are also some methods that stand independently from the above classifications that are discussed here collectively. Fast Face-swap [212] views the identity swap task as a style transfer task, achieving its goals based on VGG-Net. However, this method has poor generalization. Some methods [193, 194] apply the Transformer architecture to face swapping tasks. Leveraging a facial encoder based on the Swin Transformer [57], TransFS [194] obtains rich facial features, enabling facial swapping in high-resolution images. ReliableSwap [213] enhances the model’s identity preservation capabilities by constructing a reliable supervisor called the "cyclic triplet." However, it has limitations in preserving attribute information.
@@
Yang et al[229] AAAI’24 Exp, Pose ➊➌➍ The facial tri-plane is represented by canonical tri-plane, identity deformation, and motion components, achieving face reenactment without the need for 3D parameter model priors.
@@
Bounareli et al[232] IJCV’24 Exp, Pose ➊➌ In GANs’ latent space, head pose and expression changes are decoupled, achieving near-real outputs through real image embedding.
@@ 3.1.2 Face Reenactment
$\bullet$ Feature Decoupling. The latent feature decoupling and driving methods [16, 231, 230, 232, 229] aims to disentangle facial features in the latent space of the driving video, replacing or mapping the corresponding latent information to achieve high-fidelity facial reproduction under specific conditions. HyperReenact [16] uses attribute decoupling, employing a hyper-network to refine source identity features and modify facial poses. StyleMask [231] separates facial pose and expression from the identity information of the source image by learning masks and blending corresponding channels in the pre-trained style space S of StyleGAN2. HiDe-NeRF [230] employs a deformable neural radiance field to represent a 3D scene, with a lightweight deformation module explicitly decoupling facial pose and expression attributes.
$\bullet$ Self-supervised Learning. Self-supervised learning employs supervisory signals inferred from the intrinsic structure of the data, reducing the reliance on external data labels [233, 236, 234, 237]. Oorloff et al[234] employs self-supervised methods to train an encoder, disentangling identity and facial attribute information of portrait images within the pre-defined latent space itself of a pre-trained StyleGAN2. Zhang et al[237] utilizes 3DMM to provide geometric guidance, employs pre-computed optical flow to guide motion field estimation, and relies on pre-computed occlusion maps to guide the perception and repair of occluded areas.
TABLE IV: Overview of representative talking face generation methods. Notations: ➊ LRW, ➋ VoxCeleb2, ➌ MEAD, ➍ Self-build, ➎ LRS2, ➏ HDTF, ➐ LRS3, ➑ CREMA-D, ➒ VoxCeleb, ➓ FFHQ. Full table can be viewed in Appendix B.
@@
MakeItTalk [238] TOG’20 Uable to control pose and emotional variations well. Separating content information and identity information from audio signals, combining LSTM and self -attention mechanism to enhance head movement coherence.
Ji et al[239] CVPR’21 ➊➌ Inability to control pose and emotional variations. By breaking down the input audio sample into content and emotion embeddings, cross-reconstruction of emotional disentanglement creates facial landmarks with nuanced emotional content.
@@
EmoTalk [242] ICCV’23 ➍➏ Poor real-time performance and expression details. An emotion-entangled encoder and emotion-guided decoder enable emotion injection, with outputs generated using Blendshape and FLAME model rendering.
@@
Wang et al[249] TPAMI’24 ➋➌➏ Unable to delicately control emotions. Using 3DMM as an intermediate variable to convey facial expressions and head movements, and intro -ducing additional reference videos to extract the desired speaking style.
@@
3D-Model AD-NeRF [254] ICCV’21 Inadequate control of emotions and latent attributes. The NeRF based approach achieves accurate reproduction of detailed facial components and generates the upper body region.
Ye et al[258] ICLR’24 [76] Lack emotional control and occasional artifacts. Facial and audio information is separately represented using tri-plane, followed by rendering. The gen -erated results are further optimized based on the super-resolution network.